I have to solve the system of equationg
7x+5y=10
-2x+2y=4
Please help, thank you :)

Answer by stanbon(57250)

solve the system of equation:
7x+5y=10
-2x+2y=4
------
Multiply thru 1st by 2.
Multiply thru 2nd by 7.
-------------------
14x + 10y = 20
-14x + 14y = 28
--------------------
Add and solve for "y":
24y = 48
y = 2
----
Solve for "x":
7x + 5y = 10
7x + 10 = 10
x = 0
================
Cheers,
Stan H.
================